Bread and Butter are the two complementary goods that people often prefer to consume for breakfast in the weekend. The utility function for the two goods are:

U (B, M) =   min (4B, 16M)

  •  Find the demand function for Bread (B) and Butter (M) using the above utility function. 
  • What would be the quantities of Bread (B) and Butter (M) if an individual is willing to spend $20 on Bread (B) which costs $2, and Butter (M) which costs $4.
1
Expert's answer
2021-09-12T19:33:35-0400

Solution:

1.). Find the demand functions:

Derive the marginal utility functions:

MUB = "\frac{\partial U} {\partial B} = \frac{64} {B}"


MUM = "\frac{\partial U} {\partial M} = \frac{64} {M}"


"\frac{MU_{B} }{P_{B}} =\frac{MU_{M} }{P_{M}}"

PB = 2

PM = 4

Solve for B:

"\frac{64}{B} \div 2 = \frac{64}{M} \div 4"

B = 2M

The demand function for Bread (B): B = 2M

Solve for M:

"\frac{64}{B} \div 2 = \frac{64}{M} \div 4"


"M = \frac{B}{2}"

The demand function for Butter (M): M = "\frac{B}{2}"

 

2.). Find the quantities of Bread (B) and Butter (M) if an individual is willing to spend $20:

Derive the budget constraint:

I = PbB + PmM

Where: I = Income = 20

           Pb = Price of B = 2

           Pm = Price of B = 2

20 = 2B + 4M

B = 2M

Substitute in the budget constraint:

20 = 2(2M) + 4M

20 = 4M + 4M

20 = 8M

M = 2.5

B = 2M = 2 "\times"2.5 = 5

UBM = (5, 2.5)

The quantities of Bread (B) and Butter (M) if an individual is willing to spend $20 are 5 and 2.5